OpenStudy (anonymous):

Find the quotient: (16x^3+24x^2)/2x

OpenStudy (anonymous):

\(\Large \frac{16x^3+24x^2}{2x}=\frac{16x^3}{2x}+\frac{24x^2}{2x}=8x^2+12x\)